### v1

# A Neural Rewriting System to Solve Algorithmic Problems

Abstract
Modern neural network architectures still struggle to learn algorithmic procedures that require to systematically apply compositional rules to solve out-of-distribution problem instances. In this work, we propose an original approach to learn algorithmic tasks inspired by rewriting systems, a classic framework in symbolic artificial intelligence. We show that a rewriting system can be implemented as a neural architecture composed by specialized modules: the Selector identifies the target sub-expression to process, the Solver simplifies the sub-expression by computing the corresponding result, and the Combiner produces a new version of the original expression by replacing the sub-expression with the solution provided. We evaluate our model on three types of algorithmic tasks that require simplifying symbolic formulas involving lists, arithmetic, and algebraic expressions. We test the extrapolation capabilities of the proposed architecture using formulas involving a higher number of operands and nesting levels than those seen during training, and we benchmark its performance against the Neural Data Router, a recent model specialized for systematic generalization, and a state-of-the-art large language model (GPT-4) probed with advanced prompting strategies.
1 Introduction
Whether neural networks can learn to perform systematic reasoning is a long-standing problem in artificial intelligence and cognitive science Fodor and Pylyshyn (1988), and the recent success of deep learning in solving challenging reasoning tasks has reinvigorated this debate Hupkes et al. (2020). However, despite some promising achievements Lake and Baroni (2023), there is general consensus that even state-of-the-art transformer architectures and large language models still lack systematic and compositional generalization skills Csordás et al. (2022a); Press et al. (2023); Dziri et al. (2023).
In this work, we investigate this issue by considering a class of reasoning problems that requires to simplify symbolic formulas to a minimal form. Despite its apparent simplicity (these problems could be effectively tackled by automatic symbolic solvers), learning to solve this kind of problems is extremely challenging, as it implies the discovery of algorithmic procedures that generalize to formulas with different nesting depths and number of operands.
Inspired by rewriting systems from the tradition of symbolic artificial intelligence Dershowitz and Jouannaud (1990) and by the evidence that learning algebra retrains our visual system to perceive the hierarchical structure of symbolic expressions Marghetis et al. (2016), we propose a neural architecture that decomposes formula simplification problems into sub-tasks carried out by specialized neural modules.
We compare our approach to the Neural Data Router Csordás et al. (2022b), an end-to-end trained variant of a transformer encoder designed to solve algorithmic problems, and to GPT-4 OpenAI (2023) using self-consistency zero-shot CoT prompting Wang et al. (2023), which improves performance of large language models on reasoning tasks. We show that our approach is more effective than the alternatives, especially when solving deeply nested formulas. We also analyze the causes of failure of our system: we find that accurately selecting the portion to be simplified in very long formulas is the hardest sub-task, shedding light on the limitations of current approaches to deal with length generalization in algorithmic tasks with transformers.
2 Related works
Systematic Generalization A consistent stream of research investigates the extent to which neural networks can learn to reason in a systematic and compositional way Hupkes et al. (2020); Testolin (2024). Several benchmarks have been proposed to measure this capacity: among the most popular, ListOps Nangia and Bowman (2018) and CTL Liska et al. (2018) can be seen as instances of formula simplification problems, in which the learning system is tasked to produce the evaluation of a formula which typically requires to compute intermediate results, either implicitly or explicitly. Other formula simplification benchmarks include derivation and integration problems Lample and Charton (2019) or polynomial simplification Agarwal et al. (2021).
Modern neural architectures achieve different degrees of systematic compositional generalization Ontañón et al. (2022); Csordás et al. (2022b); Lake and Baroni (2023). Interestingly, the emergence of specialized modules in Mixture of Experts architectures has been recently studied by Mittal et al. (2022), who found that while modular architectures can outperform monolithic ones in both in- and out-of-distribution testing conditions, highly specialized modules do not emerge from standard end-to-end training without ad-hoc learning biases.
Rewriting systems and neural networks The idea of implementing neural architectures inspired by symbolic rewriting systems was initially explored in unsupervised learning settings, where individual network weights represented string tokens to be rewritten Komendantskaya. (2009). Others implemented a rewriting system for algebraic problems using custom feature engineering methods and a feed-forward neural network Cai et al. (2018), while in Chen and Tian (2019) the authors propose a reinforcement learning-based system that can learn a general rewriting mechanism by first choosing a region to simplify and then the appropriate rewriting rule to apply to the chosen region.
LLMs Large Language Models have recently been tested on complex reasoning tasks, often exhibiting impressive problem solving capabilities. These tasks often include ‘symbolic reasoning tasks’, such as coin flip, last letter concatenation Wei et al. (2022) and boolean variable assignment Anil et al. (2022), which can be solved by applying simple algorithms. However, LLMs still struggle with problems that require systematic generalization Press et al. (2023).
3 Formula simplification problems
Many compositional generalization benchmarks require to evaluate a formula which can be made arbitrarily long by replacing elementary operands with expressions that have an equivalent value. For example, in ListOps each integer operand can be replaced with an operation on a list of integers which evaluates to that operand.
In this class of problems, formulas $f\in F$ are composed of two semantically distinct elements: operators $O$ and arguments [^1] We also sometimes refer to arguments as operands. We prefer to use the term arguments here to avoid confusion with operators $O$. $A$. Arguments can be either atomic elements $E$, such as integers, which are also the final values of any formula, or other formulas. Therefore, in general formulas have a recursive structure (see Figure 1). Furthermore, since in these problems each formula has always one and only one final value, any formula has the generic compact form $f=o(a_{1},...,a_{n})=e$, where $o\in O$, $e\in E$ and $a_{j}\in F\cup E,~{}\forall j\in[1,n]$. We can define leaf formulas $F^{L}\subset F$ as the subset of all formulas with form $f^{L}=o(a_{1},...,a_{n})\;\ \mathrm{s.t.}\;\ a_{k}\in E,~{}\forall k\in[1,n]$.
Considering a formula as a sequence of symbols $f=f_{1},f_{2},...,f_{N}$, we can identify the sub-sequences $f^{L}=f_{j},f_{j+1},...,f_{k},j<k$ which correspond to leaf formulas. For any problem, we can further define the set of rewriting rules $r\in R$ which map leaf formulas to their values: $r:F^{L}\rightarrow E,~{}\forall r\in R$. In these problems, some assumptions are made about rewriting rules in $R$. First, rules are assumed to apply to contiguous sub-sequences of symbols. Second, each rewriting rule is assumed to be contractive, that is to map sub-sequences to shorter ones. Third, the rewriting rules have mutually exclusive applicability scope, i.e. given a leaf formula, one and only one rule can be applied to reduce it. Fourth, several leaf formulas could be present in a formula at once, but the order in which rewriting rules are applied on them is assumed to be irrelevant.
According to the formalization outlined above, any algorithmic task which falls into this class of problems can be solved by iterative rewriting the initial formula applying a finite sequence of valid rewriting rules. In general, if at any point in the iterative process there is no valid rule that can be applied to a leaf formula, the computation ends. More precisely, we can identify three sub-tasks which should be performed to solve these problems: selecting a sequence of symbols which corresponds to a leaf formula; solving the leaf formula applying an appropriate rewriting rule; and replacing the leaf formula with the output of the rewriting rule in the original formula.
The task we are interested in this paper consists in learning from a set of (short) formulas with reductions =$\mathcal{F}=\{(f_{i},e_{i})\}_{i=1}^{d}$, where $e_{i}\in E$ is the reduction of formula $f_{i}$, a rewriting system capable of strong out-of-distribution generalization, i.e. able to correctly reduce formulas that are much more complex of the ones in the training set. In addition, we are interested in understanding how far a structured neural architecture can be successful in completing that task.
4 Neural Rewriting System
Figure 1: Examples of solution of ListOps, Arithmetic and Algebra formulas. Leaf formulas are highlighted in yellow.
Figure 2: Schematic representation of the Neural Rewriting System.
In this section, we describe the proposed Neural Rewriting System, a neural architecture composed of three integrated modules, i.e., Selector, Solver, Combiner, where each module is responsible for the solution of a sub-task required to solve a formula by iterative rewriting. A schematic representation of the system is shown in Figure 2.
4.1 Selector Module
In analogy to what happens in humans when they deploy object-based attention to locate algebraic sub-expressions that can be simplified Marghetis et al. (2016), the Selector is trained to identify the last leaf formula [^2] To enable generalization on formulas with more arguments than seen during training, if the identified leaf formula has more than two arguments, we train the Selector to output the smallest fraction of the leaf formula which can be simplified. This corresponds to the operator and the two first arguments appearing in the leaf formula. occurring in the input formula on which a rewriting rule can be applied. When training the Selector, we assume that it will always receive syntactically correct formulas, i.e. we use teacher forcing. The identified leaf formula is then given in input to the Solver. At inference time, the output of the Selector is also used to dynamically set the filters of a Convolutional Neural Network that is a component of the Combiner. Therefore, this module can also be seen as a HyperNetwork Ha et al. (2017) which sets the weights of the Combiner with discrete values.
Since we represent problem instances as strings of text, we frame the problem as a sequence-to-sequence task and use a variant of the transformer Vaswani et al. (2017) to implement the Selector. We modify the vanilla transformer in two ways. First, given the class of problems we consider in this work, we can assume that the problem the Selector needs to solve has a local nature: the leaf formula to be simplified is always located in a limited region of the input sequence. We therefore mask all entries in the self-attention matrix of the encoder but the ones around the main diagonal (i.e., we make them $-\mathrm{inf}$). The diagonal window is $12k+$ 1 tokens wide, where $k$ is an hyperparameter. Second, to improve generalization to out-of-distribution problem instances we require the Selector to be able to identify a leaf formula independently of the length of the input sequence, provided that it is syntactically correct. Following recent evidence that the ability of transformers to generalize on longer samples can be influenced by the choice of positional encodings Ruoss et al. (2023); Csordás et al. (2021), we use Label-based Positional Encodings Li and McClelland (2022) to improve the capability of the Selector to identify leaf formulas in very long sequences. The positional information of an input sequence of $L$ tokens is thus encoded in the following way: given a sequence of $N$ sinusoidal positional encodings, where $N$ is a large number that represents the maximum expected length of an input, $L$ integers are sampled in the interval $[0,N-1]$ and then sorted. The encodings found in the positions corresponding to the sampled integers are then summed to the embeddings of the tokens in the input sequence.
Multi-output generation To further increase the accuracy of the Selector on longer sequences, we employ a multi-output generation strategy. We produce several output sequences sampling from the probability distribution derived from the decoder’s outputs and then we choose the best one taking into account both a measure of confidence of the Selector and a measure of input-output agreement computed by the Combiner.
Specifically, we can see each output of the Selector as a candidate leaf formula $\hat{f}^{L}$. We generate any token $\hat{f}_{i}^{L}$ in an output sequence $\hat{f}^{L}$ by sampling from the probability distribution obtained applying the $\mathrm{softmax}$ function to the logits produced by the final fully-connected layer of the decoder. We do not use any temperature parameter when sampling the output tokens. For any input formula $f$, we repeat the stochastic generation process $M$ times, thus generating a sequence of candidate leaf formulas $\hat{F}^{L}=\langle\hat{f}^{L,1},...,\hat{f}^{L,M}\rangle$. We define the confidence of the Selector on any $\hat{f}^{L,j},1\leq j\leq M$ as the joint probability of sampling its tokens: $c(\hat{f}^{L,j})=\prod_{i=1}^{N}p_{i}^{j}$, where $N$ is the number of tokens in $\hat{f}^{L,j}$, and $p_{i}^{j}$ is the probability to sample token $\hat{f}_{i}^{L}$ in $\hat{f}^{L,j}$. We also define an agreement score $a(\hat{f}^{L,j},f)\in[0,1]$ which gives information about the fraction of $\hat{f}^{L,j}$ which is exactly present in the input formula $f$. This measure is computed by the Combiner and thus it is formally defined in Section 4.2. We then select the final output $f^{L}$ of the Selector as the one with the highest Selector confidence which has an agreement score equal to $1$ — that is, it matches the input sequence exactly. More formally, $f^{L}=\hat{f}^{L,j}\in\hat{F}^{L}\;$ $\mathrm{s.t.}\;\ c(\hat{f}^{L,j})\geq c(\hat{f}^{L,k})~{}\forall j,k\in[1,M]\land a(\hat{f}^{L,j},f)=1$.
Dynamic windowing We implement a dynamic windowing mechanism on longer input sequences that allows us to increase the model’s generalization capacity on complex problem instances. We apply this mechanism on top of the multi-output generation strategy by modifying its behavior for sequences longer than a given threshold $T$. Given an input formula $f$, if $|f|<T$ the computation is executed as described in the previous paragraph. Otherwise, we define a sequence of $M$ copies of the input $\langle f^{(1)},...,f^{(M)}\rangle$ corresponding to the $M$ candidate leaf formulas $\langle\hat{f}^{L,1},...,\hat{f}^{L,M}\rangle$ generated by the Selector. The length of each $f^{(k)}$ will be reduced by applying a window function $w$. Considering any input $f$ as a sequence $f_{1},...,f_{N}$ of $N$ tokens, we define the window function $w(f,k)=f_{k+1},...,f_{N}$ which reduces the length of the input by giving as output its last $k$ tokens. Since the Selector is trained to output the last leaf formula appearing in the input, the window function reduces the input length starting from the first tokens. We divide the sequence of copies of the input $\langle f^{(1)},...,f^{(M)}\rangle$ into 10 groups $F^{(1)},...,F^{(10)}$ of equal size. Intuitively, in each group the length of the input is reduced by a different percentage of tokens. More formally, the window function will be parameterized by $k=\mathrm{floor}(|f^{(i)}|\cdot\frac{j}{10})~{}\forall f^{(i)}\in F^{(j)},~{}\forall j\in[1,10]$. This ensures that the model can observe the whole input sequence and also select a leaf expression in the portion of the sequence where it can identify one with highest confidence.
4.2 Solver Module
Differently from classical rewriting systems, we do not represent rewriting rules symbolically in a dictionary of rules. Instead, we aim to learn rewriting rules from data using a neural network. Given a leaf formula $f^{L}$ generated by the Selector, the Solver is trained to output the corresponding reduction $e$, thereby implementing an appropriate rewriting rule for the formula. When training the Solver, we assume that the Selector will always produce perfect outputs (teacher forcing). The Solver is also trained to identify the termination state of the computation. That is, when the formula has reached its final state, it is the Solver’s responsibility to signal that such state has been reached. The Solver thus also learns to map atomic elements, which correspond to the final value of a formula, to the special symbol $\omega$ that represents the end of computation.
As in the case of the Selector, we frame the task as a sequence-to-sequence task. In this case, since the Solver has to compute simple mappings without any particular assumption, we implement it using a vanilla transformer.
4.3 Combiner Module
The last module in the architecture is the Combiner, whose goal is to produce a simplified version of the original formula, given the formula itself $f$, the leaf formula $f^{L}$ identified by the Selector and its reduction $e$ computed by the Solver.
In order to carry out its task, the first operation that the Combiner must perform is finding the position in $f$ where the leaf formula $f^{L}$ appears. We notice that the convolution is a suitable operation to detect which portion of an input sequence has the highest match with another sequence used as a filter, so we implement this operation using a 2D Convolutional Neural Network (CNN) whose filters are set dynamically at execution time using the output of the Selector, rather than being learned with backpropagation.
More precisely, we represent both the input sequence $f$ and the leaf formula $f^{L}$ as sequences of 1-hot vectors over the same vocabulary. Since the leaf formulas found for different sequences in a batch can have different lengths, we pad each one with zeros to prevent the padding to match in the input. Then, we set the filter of the 2D CNN to the 1-hot representation of $f^{L}$. We refer to the CNN parameterized in this way as $\mathrm{CNN}_{f^{L}}$. Doing so allows us to obtain from the output of the convolution both information on the location of the best match of $f^{L}$ in $f$ and on the number of tokens in $f^{L}$ that match $f$ exactly in some point. Indeed, we can compute the location of the best match as $\mathrm{pos}(f^{L},f)=\mathrm{argmax}~{}\mathrm{CNN}_{f^{L}}(f)$. Furthermore, we can compute the agreement score $a(f^{L},f)=\frac{\mathrm{max}~{}\mathrm{CNN}_{f^{L}}(f)}{|f^{L}|}$, where $|f^{L}|$ is the number of tokens in $f^{L}$. Dividing by $|f^{L}|$ makes the score normalized, which allows us to compare the agreement scores of leaf formulas with different lengths. Indeed, as described in Section 4, the Selector uses this score in the multi-output generation strategy to discard the outputs that do not have an exact match in the input formula. Notice that in this case the CNN is parameterized using candidate leaf formulas $\hat{f}^{L}_{j}$ whose accuracy scores with $f$ are compared. If there is no Selector output such that $a(\hat{f}^{L}_{j},f)=1$, the computation on the input sequence $f$ is interrupted, and this is considered as a failure of the model.
After finding the position of the leaf expression in $f$, the Combiner replaces $f^{L}$ with $e$ in $f$, to compute the simplified version of the formula $f^{\prime}$. We implement this operation as a deterministic operator with input $f$, $f^{L}$, $e$, and $\mathrm{pos}(f^{L},f)$.
5 Experiments and results
(a) ListOps
(b) Arithmetic
(c) Algebra
Figure 3: Performance of the Neural Data Router, GPT-4 and Neural Rewriting System on the ListOps, Arithmetic and Algebra problems.
We evaluated the capacity of our model to solve out-of-distribution problem instances by considering three cases of formula simplification tasks: ListOps, arithmetic, and algebraic formulas. We benchmark our Neural Rewriting System against the Neural Data Router and GPT-4, which is one of the most advanced LLMs currently available. Code and data used in our experiments are available as Supplementary Materials and will be made public upon acceptance.
5.1 Datasets
For all three tasks, the automatic generation of formulas is parameterized by specifying the nesting level of a formula, that is the maximum number of nested formulas that can be found as arguments of the top-level formula plus the level of the formula itself. At each level, any formula is then nested in two points: exactly two arguments in the formulas on that level will be other formulas.
ListOps The ListOps dataset Nangia and Bowman (2018) was introduced to evaluate the capacity of neural networks to build parse trees of nested formulas. The original dataset consists of formulas composed of operations on lists of integers, including minimum, maximum, median and sum modulo $10$ of a list of integers. We modified the ListOps dataset to generate expression that comply with the schema described above. In particular, we fixed the number of nesting points at each level of a formula to 2, and we made it possible to specify the number of arguments that appear in formulas at any level. To reduce the complexity of the problem, we reduced the set of operations to minimum, maximum and sum modulo $10$.
Arithmetic We generated formulas composed by sum, subtraction and multiplication operations between two integers sampled in the interval $[-99,99]$. Since in this work we do not investigate the capacity to generalize to numbers with more digits than those seen during training, we used the modulo $100$ of the intermediate results in the solution process.
Algebra We considered a subset of algebraic formulas which can always be reduced up to a minimal form in a deterministic way. Such formulas involve sum and subtractions between two monomials, and their final value is always a monomial. The numerical coefficients of monomials were sampled in the interval $[-99,99]$, and each monomial can contain up to four literal variables sampled from the set $\{a,b,x,y\}$. All monomials appearing in a given formula have the same literal variables. As in the case of Arithmetic, all intermediate numerical values were taken modulo $100$ when computing the final value of the formula.
5.2 Experiments
5.2.1 Neural Rewriting System
We describe here how we built the training and validation sets for the Selector and Solver modules. We provide the details of the training procedure and hyperparameter selection for the models in the Technical Appendix.
Selector We have included in the training set for the Selector both top-level formulas and formulas that appear as intermediate solution steps in their simplification process. We did so in order to show the Selector during training the complete solution process of formulas up to a some level of difficulty. Indeed, we aim to learn a selection mechanism that can be applied also to more complex formulas.
In Arithmetic and Algebra problems we included in the training set formulas with nesting levels 1, 2 and 3, together with formulas that represent intermediate steps in the resolution process of these formulas. In the ListOps problem, in which we could specify the number of arguments that appear in the formulas, we included in the training set formulas with nesting levels 1 and 2 that have 2 or 3 arguments for each operator, together with formulas appearing in their resolution process. In all three problems, the training set also contains atomic elements which correspond to the final value of the initial formula. During training, we generated batches containing, with equal probability, samples with different structural complexity.
We constructed a separate in-distribution validation set with samples having the same structural characteristics as the those in training set. Since the Selector should possess out-of-distribution (OOD) generalization capabilities to identify the leaf formulas even in longer inputs than those seen during training, we also built a separate out-of-distribution validation set containing formulas with greater structural complexity. As done in the training set, we also included formulas representing examples of intermediate resolution steps. For Arithmetic and Algebra, we included samples with nesting levels 4, 5 and 6, together with formulas that appear in their resolution process. For ListOps, we included in the OOD validation set formulas with nesting levels 3 and 4 that have 2, 3 or 4 arguments for each operator, together with formulas that appear in their resolution process. In order to control the structural complexity of the formulas, the OOD validation sets are balanced across nesting level of the leaf expressions [^3] In the case of ListOps we also balance across the number of arguments in the leaf expression..
Solver The training and validation set of the Solver are composed by two kinds of samples: atomic elements and leaf formulas. During training, we generated batches including both kinds of samples with equal probability.
5.2.2 Neural Data Router
The Neural Data Router (NDR) is a variant of a transformer encoder designed to learn to solve algorithmic problems with strong out-of-distribution generalization capacity. This model has been previously evaluated on relatively simple algorithmic benchmarks that are nevertheless significantly similar to the formula simplification problems we consider, including the solution of formulas in the original ListOps dataset and simple arithmetic formulas on single-digit integers. The main differences with the problems we consider here is that the formulas that appear in our problems can have in general many leaf expressions and therefore can be simplified in more than one point, and the operands in our Arithmetic and Algebra problems have a greater complexity. Moreover, in the ListOps problems we aim to generalize on the number of operands as well as on the nesting depth of the formulas, whereas in the original work the number of operands in each expression was fixed. Furthermore, we use significantly less samples during training, and of lower complexity, as the NDR was originally trained on ListOps formulas with up to 5 nested operations and a variable number of operands between 2 and 5.
The problems considered in the original work always had a single-digit integer as a result, which the model was trained to output as the first token in the sequence given as output by the encoder. Since this no longer apply in our problems, except in the case of ListOps, in general we read the final answer from the first $k$ positions of the sequence produced by the encoder, where $k$ is the maximum length of a problem’s targets.
We build all development sets for the NDR using the same top-level formulas we included in the analogous sets for the Selector. Since the NDR was designed to compute intermediate results sub-symbolically, we use as training target the final simplified form of the input formula and we do not include intermediate formulas as we do with the Neural Rewriting System. Following the original experimental protocol, we made the training set balanced across nesting level and number of operands. As done when training the Selector module, we create in- and out-of-distribution validation sets to optimize hyperparameters, using the Weights and Biases Bayesian search on the same hyperparameters and intervals described in the original work. We report the final hyperparameter values for each task in the Technical Appendix.
5.2.3 GPT-4
Research on prompt engineering techniques has shown that appropriate prompting method can boost the reasoning capabilities of LLMs. We choose to prompt GPT-4 using self-consistency prompting Wang et al. (2023) combined with zero-shot CoT Kojima et al. (2022).
Zero-shot CoT has been introduced as a simpler alternative to Chain-of-Thought (CoT) prompting, which allows to achieve similar performance on reasoning benchmarks without the need to engineer exemplars for few-shot reasoning. This is achieved by simply making the answer of the model start with the sentence: “Let’s think step-by-step”. Self-consistency prompting is based on the idea that reasoning problems can often be solved following different reasoning paths that can lead to the same conclusion. Therefore, letting the model generate several outputs and then choosing the one which was generated more consistently can increase the confidence in generated output and improve accuracy. To this aim, we generate 10 outputs for each input, and indeed observe a significant improvement in the model’s performance. We report the performance of GPT-4 with zero-shot CoT
The actual prompt used for zero-shot CoT was built giving the model a minimal description of the problem at hand and then asking to solve it. For example, the zero-shot CoT prompt corresponding to the ListOps input sample [MIN[SM54][MIN39]] is: “MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN [SM 5 4] [MIN 3 9]].”. Following the zero-shot CoT prompting technique, the model was then prompted a second time to extract the well-formatted final answer.
5.3 Results
We report the performance of all models considered in Figure 3. All models are tested on 100 samples for each level of complexity. We measure the sequence accuracy of the models, that is the exact match between the model’s output and the target sequence. When we evaluate GPT-4 on the Algebra task, we also take into account equivalent forms of the symbolic outputs using the SymPy library Meurer et al. (2017).
We first observe that all the models have a better performance on simple data splits, which tends to degrade on more complex splits. This indicates that increasing the number of arguments and the nesting of formulas makes the problems more difficult for all models, as expected.
Analysing the performance of the Neural Data Router, we find that the models with the best hyperparameters configurations have a lower accuracy than other models on the simple splits and are not able to generalize the learned solution process to complex problem instances.
We observe that GPT4 can achieve perfect or almost perfect results on the simpler problem instances, surpassing the performance of both the Neural Data Router and our modular architecture. However, the performance of GPT-4 is significantly lower on more complex instances of the problems. Indeed, while self-consistency prompting improves the model’s performance compared to vanilla zero-shot CoT prompting [^4] We report the performance of GPT-4 with vanilla zero-shot CoT prompting in the Technical Appendix for comparison., it has a limited impact on the capacity of the LLM to solve complex formulas. Comparing the performance achieved on ListOps to that on the other two tasks, we can conjecture that solving deeply nested formulas is harder for GPT-4 than solving formulas with many operands. Furthermore, it seems that accurately computing many arithmetic operations between double-digit integers, including multiplication modulo 100, is the hardest task for the LLM.
Figure 4: Errors committed by the Neural Rewriting System classified by cause of error. Error bars corresponding to a complexity split refer to models where the Selector generates 10, 100 or 1000 outputs per input, and employs the dynamic windowing mechanism. Complexity splits are defined by their nesting level N and number of arguments A.
Figure 5: Input length against the average confidence score of 1000 outputs of the Selector. The vertical lines represent the maximum input length in the training set.
As for the Neural Rewriting System, generating 1000 outputs per input in the Selector allowed to achieve a higher degree of out-of-distribution generalization on more complex samples than those seen by the Selector during training. The performance of the model is consistently equal or superior to that of the Neural Data Router on both simple and complex data splits. Compared to GPT-4, the model achieves a similar performance on the simplest data splits, but its accuracy of is higher than that of GPT-4 on data splits of medium difficulty, such as arithmetic expressions of nesting level 4 and 5 and algebraic expression of nesting level 3 and 4. However, also the accuracy of the Neural Rewriting System drops dramatically on the most difficult arithmetic and algebraic formulas.
We analysed the root cause of the errors committed by the NRS model for each difficulty split in Figure 4. We also assessed the impact of the multi-output generation strategy on the model’s performance by comparing the errors committed by the model when 10, 100 and 1000 outputs per input were generated by the Selector. We classify the cause of the errors into three categories: errors caused by the Selector when it identifies a leaf formula which is not present in the input formula; errors caused by the Selector when it identifies a leaf formula that is present in the input formula but does not correspond to a well-formed expression and whose substitution thus causes the corruption of the input formula; and errors caused by the Solver when it outputs a wrong solution for the leaf formula identified by the Selector. We call these selector/missing, selector/corrupted and solver errors, respectively.
Looking at the frequency of each of these classes of errors, we clearly see that the main cause of error made on the hardest split is the Selector, either due to non-existing leaf formulas or ill-formed ones. However, we can also recognize that the multi-output generation strategy is effective for simpler data splits: generating many Selector outputs effectively reduces the amount of errors it causes, and especially the ones due to the identification of a non-existing leaf formula.
Since the confidence score plays an important role in the selection of the best Selector output in the multi-output generation strategy, we analysed the variation of the score corresponding to variations of length of the input sequences in the test set. Figure 5 represents input lengths versus the average confidence score of 1000 Selector outputs. The vertical lines in the plot represent the maximum input length seen during training. While the average confidence score remains close to the training range for outputs that are moderately longer than the longest training samples, it drops significantly for sequences much longer than that. We also notice that confidence scores are lower for very short arithmetic and algebraic formulas. We suppose this is due to the fact that the number of formulas of a given nesting level grows exponentially with the value of nesting level. Therefore, the Selector observed fewer examples of formulas with nesting level 1 during training. However, this does not hinder performance when the Selector generates many outputs per input.
The dynamic windowing mechanism effectively addresses the drop of confidence on longer sequences. We report in Figure 3 the accuracy of a Neural Rewriting System in which the Selector generates 1000 outputs per input and applies the dynamic windowing mechanism on the longest sequences. We selected the length threshold $T$ that is used to specify which sequences will be processed using dynamic windowing based on the analysis of the confidence scores of the Selector discussed above. Specifically, we set the thresholds to 60, 150 and 200 for the ListOps, Arithmetic, and Algebra problems, respectively. The analysis of errors in Figure 4 shows that this mechanism improves the model’s performance by further reducing the number of errors caused by the Selector on the most complex data splits in all problems.
6 Discussion
In this work, we proposed an original approach to solve formula simplification problems which we dubbed Neural Rewriting System. We described how such problems can be generally solved iteratively by first identifying a leaf formula, then computing its value, and finally replacing the value in the original formula. We show that iteratively decomposing in such a way the solution process and implementing a specialized module for each sub-task, it is possible to achieve higher accuracy on the solution of complex formulas than both a highly-specialized neural architecture trained end-to-end and a general purpose Large Language Model.
Limitations While we consider formula simplification problems with different number and kinds of operands, we also pose several constraints on these formulas. For instance, we limit our study to formulas which have only one final form which can always be reached independently of the substitution applied. Both of these conditions are not satisfied, in general, by algebraic expressions, whose form could be made arbitrarily complex and not further reducible. Future work should thus consider the more general case of formulas with several possible final forms, reachable from different solution paths. This scenario would require learning a decision mechanism to select appropriate rewriting rules, for example by using reinforcement learning. Furthermore, the architectural mechanisms that we employ to generalize on formula simplification problems rely on the exploitation of specific assumptions that apply to the class of problems consider in this study. More general architectural elements and/or training methods should thus be devised to achieve more robust and widely-applicable generalization in neural systems.